OpenStudy (anonymous):

6th grade math help!~

OpenStudy (.whitedragon.):

whats the question?

OpenStudy (anonymous):

What is the product?

OpenStudy (hippoharry):

http://www.mathpapa.com/ use MathPapa and it helped you get the answer and explain it

OpenStudy (abdullahm):

Do the multiplication like normal ignoring the signs. And then look at the signs. A negative times a negative becomes a positive. A positive times a negative becomes a negative.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

is it 105?

OpenStudy (.whitedragon.):

-105

OpenStudy (.whitedragon.):

use mathway https://www.mathway.com/

OpenStudy (anonymous):

but @AbdullahM said - x - = + :\

OpenStudy (.whitedragon.):

-105

OpenStudy (shadowlegendx):

\[-5 \times (-3) \times (-7) \times (2) \times (0.5)\]

OpenStudy (shadowlegendx):

-5 times -3 = 15 15 times -7 = -105 Since 2 times 0.5 is 1, we don't have to do any further work. -105 is your answer

OpenStudy (shadowlegendx):

As it would just be -105 times 1 = -105

OpenStudy (anonymous):

Thanks!

OpenStudy (shadowlegendx):

No problem :)

OpenStudy (shadowlegendx):

Remember: Negative times Negative = Positive Positive times Positive = Positive Negative times Positive = Negative Positive times Negative = Negative

OpenStudy (.whitedragon.):

theres a special smiley face thing I learned in 6th grade pre ap and theres also a tick tac toe type graph you can do it on too